OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) greater than or equal to -12? A. x greater than or equal to -5 B. x less than or greater to -5 C. -5 less than or greater to x D. A number line with a closed circle on negative 5 and shading to the right. answer please :)

OpenStudy (texaschic101):

2(x - 1) >= -12 2x - 2 >= -12 2x >= -12 + 2 2x >= - 10 x >= -5 B is incorrect

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Thank you :) i have more if you dont mind i just neeed the answer

OpenStudy (texaschic101):

lets see them

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 3x + 1 > 10 and 4x less than or greater to 20 A. A number line with an open circle on 3, a closed circle on 5, and shading in between. B. A number line with an open circle on 3, shading to the left, and a closed circle on 5, shading to the right. C. A number line with a closed circle on 3, shading to the left, and an open circle on 5, and shading to the right. D. A number line with a closed circle on 3, an open circle on 5, and shading in between. and hers another Which of the following ordered pairs is not a solution to the inequality y less than or greater to -3x - 5? A. (-4, 0) B. (0, 0) C. (-3, 0) D. (-2, 0) i have morre but ill let u answer these before i write more

OpenStudy (texaschic101):

3x + 1 > 10 4x <= 20 3x > 10 - 1 x <= 5 3x > 9 x > 3 A is correct y < = -3x - 5 or y >= -3x - 5 I am having a problem with this one.....is it less then or equal or greater then or equal ?

OpenStudy (texaschic101):

or is their even an equal sign ?

OpenStudy (anonymous):

less than or equal to

OpenStudy (texaschic101):

give me a few minutes on this one

OpenStudy (anonymous):

Ok no problem

OpenStudy (texaschic101):

the answer is (0,0) that is NOT a solution...sorry about that

OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y <= -3x-5 does this make more sence?

OpenStudy (texaschic101):

the answer is (0,0)....I just made a mistake....this is your answer

OpenStudy (anonymous):

Ohh okay thank you so much how bout this one Solve the inequality 4x + 5 < 2x - 11 A. x > 8 B.x > -8 C. x < 8 D. x < -8 Which of the following is not an equivalent form of the compound inequality x - 1 < -12 or x + 12 > 10 x < -11 or x > -2 A number line with an open circle on -11, shading to the left, and an open circle on -2, shading to the right. A number line with an open circle on -11, an open circle on -2, and shading in between. x > -2 or x < -11

OpenStudy (texaschic101):

4x + 5 < 2x - 11 4x - 2x < -11 - 5 2x < -16 x < -8 (answer D) ======================= x - 1 < -12 x + 12 > 10 x < -12 + 1 x > 10 - 12 x < -11 x > -2 3rd one is incorrect

OpenStudy (anonymous):

thank you would you like more lol or would u like to move on to another person and ill just post as new question lol i dont like bothering ppl

OpenStudy (texaschic101):

I would help you more if I could but I am fixing to have to go somewhere. I do hope I helped you :)

OpenStudy (anonymous):

You did thank you sooo muchhh Enjoy ur dayyy

OpenStudy (texaschic101):

you too shannon...nice meeting you :)

OpenStudy (anonymous):

Youu too :)

OpenStudy (texaschic101):

c ya later :)

OpenStudy (anonymous):

Tc :)